# Is the Cure Still Worse Than the Disease? Test Overfitting by LLMs in Automated Program Repair

### v1

@@ 4.3. Overfitting with Code Refinement (RQ2)
Tests have been used to validate SWE-patches. What about using tests to improve patches at generation time? Using our test-based code refinement technique, we attempt to generate an improved code+test pair for each of the 220 code+test pairs where the tests did not initially go from fail to pass. Refinement succeeded in generating 22 pairs where the white-box test goes from fail to pass. Unfortunately, 14 out of the 22 fail black-box testing, increasing the overfitting. If we include these in our existing overfitted samples, the overfitting rate increases from 21.8% to 25.5% (Table 1). GPT-4o experiences a similar increase in overfitting, indicating that it is difficult to improve code patches by exposing tests to LLMs.
In the above experiments, the refinement loop exposed the white-box tests and additional information to the LLMs. One idea to reduce overfitting might be to hide the test and additional information, so the LLMs may not be able to trick the test. Indeed, hiding the test causes a drop in agreement between code patches and white-box tests. However, the number of instances passing black-box tests also goes down a bit. If we remove the additional information, it becomes more difficult for LLMs to overfit the patch. Unfortunately, we still see a high overfitting rate even after this intuitive overfitting reduction approach. The model can find a way by knowing that the existing test is not fail-to-pass (third row in Table 2). Note that this will not have any impact on instances that were already overfitted before the refinement.
Table 2. Test overfitting rate of repaired samples.
@@
Table 2 shows that exposing tests and prompt augmentation resolved 7 samples. Do all these 7 samples help us resolve additional instances? The answer is no. The initial code patch for one instance was already passing on the black-box test. And besides these 7, there were 2 instances that went from resolved to non-resolved. So the net gain is +4 instead of +7. That indicates the actual gain from test-based refinement is less than the apparent gain.
4.4. Impact of Revealing Black-box Tests (RQ3)
Table 3 shows how much overfitting occurs when using some of the golden tests for code selection and repair. For this experiment, we divide black-box tests into two groups: fail-to-pass reproduction tests and pass-to-pass regression tests. We only reveal the reproduction tests (so they are no longer black-box) but continue to hide the regression tests. This is a limit study, because in practice, all black-box tests are hidden. The overfitting rate, which means passing now-revealed reproduction tests but failing still-hidden regression tests, is low (13/223=5.8% and 22/194=11.3% with Claude and GPT-4o, respectively). Apart from overfitting, this also shows how well the model performs when some black-box tests are exposed. When the model passes the golden reproduction tests, the probability that the issue is resolved is high, but not 100%, because regression tests can still fail. The improvement from using black-box tests in issue resolution is modest: the number of resolved instances with the Claude model goes up from 229 to 243, which is around a 3% increase. Overfitting in this setup is much less since we are evaluating against the golden tests–however, the golden tests themselves are really just a proxy for “correctness”, against which we might still be overfitting (but we do not have a good way to measure this).
@@ 5. Discussion
5.1. LLMs’ Preference on Focal and Test
@@
5.2. Coverage for Overfitted Sample
Figure 3 compares the coverage between unbiased code patches (both white-box and black-box tests pass) and overfitted patches for Claude-3.7. Note that the tests are generated by e-Otter++ and the code patches are from Agentless. The median for unbiased patches is 1, while for overfitted patches it is less than 0.8. Also, the mean coverage is higher for unbiased patches (0.95 vs. 0.9). These results indicate that even though both unbiased and overfitted patches pass the white-box test, overfitted patches have lower coverage.
5.3. What Could Happen with Overfitting?
Frontier LLMs are capable of tricking the test and changing the code to narrowly passes that particular test. For example, instance id “astropy__astropy-14598” complains about an inconsistency in double and single quotes. The issue writer notes that while dealing with a null string, the double quotes become single quotes. To trick the white-box test, LLMs modify their initial solution (see Figure 4(a)) to a solution where “replace” is invoked three times just to pass the white-box test (see Figure 4(b)), which is clearly an indication of LLMs overfiting. We have even seen LLMs using reflection to access private methods to pass the white-box test, creating a security risk.
(a) Initial code patch
(b) Overfitted refined code patch
Figure 4. Example where test-based refinement overfits.
@@ 6. Threats to Validity
One of the main limitations of our work is that the benchmarks include only Python repositories. Therefore, our findings may not generalize to other programming languages. We experimented with two models but did not investigate model contamination or memorization. However, prior works show that model-generated tests are very different from those written by developers. On another note, our approach may be considered a way to measure contamination by exposing overfitted patches. We leave this for future research. Other minor limitations include our reliance on the coverage package for reward formulation. The Python coverage package sometimes does not work due to dependency issues or multithreading. However, this happens in less than 1% of cases, and we assign a coverage of 1 to mitigate the impact of imperfect coverage reports. Our results can also be affected by test flakiness. However, we use pre-built docker containers for the experiments, which should minimize the impact of flakiness.
@@ 7. Conclusion
8. Future Plans
This paper demonstrates that overfitting exists and shows attempts to minimize it in a test-based code repair system. However, we have not proposed any methods to address overfitting in cases where the initial white-box test changes from fail to pass on overfitted patches. This is an important problem because the majority of overfitted patches come from this category. In the future, we will try to design an approach to automatically detect these overfitted patches and propose a solution to address the issue.

### v2

@@ 4.2. Overfitting with Code Refinement (RQ2)
Tests have been used to validate SWE-patches. What about using tests to improve patches at generation time? We apply the test-based code refinement loop from Section 3 to each of the 220 instances from Figure 3 where the initial code $c_{\textrm{new}}$ failed the observed generated tests $t_{\textrm{gen}}$. This yielded 22 instances where the refined code $c^{\prime}_{\textrm{new}}$ passed the generated tests. Unfortunately, 14 out of the 22 fail the hidden tests $t_{\textrm{gold}}$, increasing the overfitting. If we include these in our existing overfitted samples, the overfitting rate increases to $\frac{50\,+\,14}{229+22}=\frac{64}{251}=25.5\%$ (see Table 1). GPT-4o experiences a similar increase in overfitting, indicating that it is difficult to improve code patches by exposing tests to LLMs.
In the above experiments, the refinement loop exposed the generated tests $t_{\textrm{gen}}$ and additional information to the LLMs. One idea to reduce overfitting might be to hide the tests and additional information, so the LLMs may not be able to trick the tests. Indeed, hiding the generated tests and revealing merely a pass/fail flag to the LLM causes a drop in agreement between code patches and generated tests. However, the number of instances passing the hidden tests $t_{\textrm{gold}}$ also goes down a bit. If we remove the additional information, it becomes more difficult for LLMs to overfit the patch. Unfortunately, we still see a high overfitting rate even after this intuitive overfitting reduction approach. The model can find a way by knowing that the existing test is not fail-to-pass (third row in Table 2). Note that this will not have any impact on instances that were already overfitted before the refinement.
Table 2. Test overfitting rate of repaired samples, and effect over hiding information to mitigate overfitting.
@@
Table 2 shows that exposing tests and prompt augmentation resolved 8 samples. Do all these 8 samples help us resolve additional instances? The answer is no. The initial code patch for one instance was already passing on the golden test. And besides these 8, there were 2 instances that went from resolved to non-resolved. So the net gain is +5 instead of +8. That indicates the actual gain from test-based refinement is less than the apparent gain.
4.3. Impact of Revealing Golden Tests (RQ3)
RQ3 asks what would happen if the test-based code refinement would not use imperfect generated tests $t_{\textrm{gen}}$ but rather golden tests $t_{\textrm{gold}}$. This is a limit study, because in practice, issues do not usually come with oracular executable acceptance tests. Table 3 shows how much overfitting occurs in this setting. Here, the final determination of whether an instance is resolved is based on passing both golden tests $t_{\textrm{gold}}$ and a subset of regression tests $t_{\textrm{old}}$. (While issue-resolution systems such as Agentless have access to the full regression testing suite contained in the original code $c_{\textrm{old}}$, they do not know which subset of those tests are expected to pass after issue resolution.) The overfitting rate, which means passing now-revealed golden tests $t_{\textrm{gold}}$ but failing the relevant subset of regression tests $t_{\textrm{old}}$, is low (13/223=5.8% and 22/194=11.3% with Claude and GPT-4o, respectively). Apart from overfitting, this also shows how well the model performs when golden tests $t_{\textrm{gold}}$ are exposed. When the model passes the golden reproduction tests, the probability that the issue is resolved is high, but not 100%, because regression tests can still fail. The improvement from using golden tests in issue resolution is modest: the number of resolved instances with the Claude model goes up from 229 to 243, around a 3% increase. This experiment emulates a situation where developers follow test-driven development and write tests prior to generating $c_{\textrm{new}}$. Unfortunately, the gains are limited, highlighting that current issue resolution systems often miss the bigger picture.
@@ 5. Discussion
LLMs’ Preference on Focal and Test
@@
Coverage for Overfitted Sample
Figure 4 compares the coverage between unbiased code patches (both $t_{\textrm{gen}}$ and $t_{\textrm{gold}}$ pass) and overfitted patches for Claude-3.7. The median for unbiased patches is 1, while for overfitted patches it is less than 0.8. Also, the mean coverage is higher for unbiased patches (0.95 vs. 0.9). These results indicate that even though both unbiased and overfitted patches pass observed tests $t_{\textrm{gen}}$, overfitted patches have lower coverage.
@@ 6. Threats to Validity
Since the benchmarks include only Python repositories, our findings may not generalize to other programming languages. We experimented with two models but did not investigate model contamination or memorization. However, prior works show that model-generated tests are very different from those written by developers. On another note, our approach may be considered a way to measure contamination by exposing overfitted patches. We leave this for future research. Other minor limitations include our reliance on the coverage package for reward formulation. The Python coverage package sometimes does not work due to dependency issues or multithreading. However, this happens in less than 1% of cases, and we assign a coverage of 1 to mitigate the impact of imperfect coverage reports. Our results can also be affected by test flakiness. However, we use pre-built docker containers for the experiments, which should minimize the impact of flakiness.